OpenStudy (anonymous):

Another Trigonometric equation. 3cot^2x-1=0

OpenStudy (campbell_st):

try 3cot^2(x) = 1 then cot^2(x) = 1/3 \[\cot(x) = 1/\sqrt{3}\] cot = 1/tan so \[\tan(x) = \sqrt{?}\]

OpenStudy (campbell_st):

? = 3 so another exact trig ratio